OpenStudy (anonymous):

Graph the equation 4x-y=3. Help anyone?? step by step would be great.

OpenStudy (shamil98):

Change the equation to y = mx + b slope-intercept form.

OpenStudy (gk99):

4x-y = 3 -y = -4x - 3 y=4x + 3 pick three points for x, then graph :)

OpenStudy (shamil98):

4x -y = 3 -4x -4x -y = -4x + 3 divide by 1 y = 4x - 3 4 is your slope (rise/run, 4/1) -3 is your y intercept at which you graph at.

OpenStudy (shamil98):

divide by -1*, typo..

OpenStudy (anonymous):

so wouldn't it be